OpenStudy (anonymous):

find the LCM y^3 +4y^2 + 4y, y^2-4y

OpenStudy (watchmath):

\(y^3+4y^2+4y=y(y^2+4y+4)=y(y+2)^2\) \(y^2-4y=y(y-4)\) So the LCM is \(y(y-4)(y+2)^2\)

OpenStudy (anonymous):

thank u